two forces 12 N and 16N are acting upon a body.what can be the maximum and minimum resultant force on the body?

Grade:12th pass

1 Answers

SAHIL
7 years ago

The resultant of two forces acting on the same point is maximum when the forces are acting in the same direction and minimum when they operate in opposite direction.

When the two forces are acting in the same direction the angle between the direction to the two forces is 0. In this case the resultant force is sum of the two forces and its direction is same as the two independent forces.

therefore net resultant maximum force=12+16= 28 N

When the two forces are acting in the opposite direction the angle between the direction to the two forces is 180 degrees. In this case the resultant force is the difference in the magnitude of two forces acting in the direction of larger of the two forces.

Thus when the forces with the given magnitude are acting in opposite direction their net magnitude will be:

IN THIS CASE MIMIMUM FORCE=16-12=4N

THANKS AND APPROVE MY ANSWER

REGARDS SAHIL